OpenStudy (anonymous):

Find the Cartesian equation of the polar equation r= 5/ (3-4sinx)

OpenStudy (anonymous):

its not simplifying out nicely -.- im gonna post what i have so far in a sec.

myininaya (myininaya):

so x is suppose to be theta right polar equations are in the from (r, theta) Cartesian equations are in the form (x,y)

myininaya (myininaya):

what you have ended up with is a hyperbola you can write it in the standard form of a hyperbola if you want

OpenStudy (anonymous):

polar coordinates is a weakness of mine =/ im using this formulas to convert: \[r^{2} = x^{2}+y^{2}, \sin(\theta) = \frac{y}{r} \rightarrow y = rsin(\theta)\]

OpenStudy (anonymous):

hey thats exactly what i needed thanks a lot again

OpenStudy (anonymous):

i dont wanna write it in standard form, im too lazy to complete the square with 40/7 >.<

myininaya (myininaya):

lol

myininaya (myininaya):

there is nothing to polar coordinates its just a bunch of substitutions and a different way of graphing

OpenStudy (anonymous):

scary stuff~

myininaya (myininaya):

i understood that some cartesian equations are easier to graph when written in polar form

OpenStudy (anonymous):

yeah, cartiods ( i think thats how you spell that), spirals, things of circular nature right?

myininaya (myininaya):

in some polar equations might be easier graphed written in cartesain form

myininaya (myininaya):

y=7 is easier to graph than rsin theta=7

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Cardioid

myininaya (myininaya):

i mean but they are the same graph but to me it is not easy to see that rsintheta=7 is a horizontal line

OpenStudy (anonymous):

yeah, its not >.< i dont see it lol